If Government Spending, Investment and Net Export are given as 250, 325 and 85 respectively, compute the equilibrium income. 


1
Expert's answer
2021-10-19T11:35:38-0400

equilibrium (Y) = consumption (c) + investment+ government purchases (g)

"y=c+i+g\n= 250+325+85= 660\ny=660"

the answer is 660